# From Intuition to Expertise: Rubric-Based Cognitive Calibration for Human Detection of LLM-Generated Korean Text

### v1

@@ B.1 Total LREAD Score Distributions
Figure 6 summarizes the distribution of total LREAD scores in Phase 2 by generator (Human, GPT-4o, Solar, Qwen2, and Llama3.1). The figure pools rater–essay evaluations across three annotators, which characterizes both central tendency and dispersion for each source. Each generator contributes six essays (two per grade level), yielding 18 pooled evaluations per generator (6 essays $\times$ 3 annotators).
While the main paper focuses on binary detection outcomes, total LREAD scores provide a graded view of rubric-guided judgments. Higher totals indicate stronger satisfaction of the rubric’s content, organization, and expression criteria, whereas lower totals reflect accumulated deficiencies revealed by the analytic protocol. Generator-wise distributions therefore provide complementary evidence about rubric-guided auditing beyond surface-level fluency. In particular, tighter dispersion is consistent with more uniform rubric outcomes across prompts and grade strata, whereas broader dispersion is consistent with greater variability in rubric-aligned writing quality.
@@
Table 6: Phase 2 total LREAD scores by grade level (mean $\pm$ SD). Total scores are summarized for each annotator and for pooled evaluator–essay evaluations. Each grade level includes 10 essays, yielding 30 pooled evaluations per level across three annotators.
@@
B.2 Annotator-Level Descriptive Statistics
Tables 6 and 7 report mean $\pm$ standard deviation (SD) of total LREAD scores in Phase 2. These summaries support two complementary interpretations. First, they report score magnitude and variability for each annotator, which contextualizes the pooled distributions in Figure 6. Second, they provide grade-level and generator-level views of total-score variation under a unified rubric.
Table 6 aggregates scores by grade level (elementary, middle, high) for each annotator and for pooled evaluations. Each grade level contains 10 essays, yielding 30 pooled evaluations per level across three annotators (10 essays $\times$ 3 annotators). These summaries contextualize rubric-guided scoring across developmental writing regimes, which differ in rhetorical maturity and normative writing proficiency.
Table 7 reports statistics by generator, enabling direct comparison between human-authored essays and each LLM source under a unified analytic protocol. Each generator contributes six essays (two per grade level), yielding 18 evaluations across three annotators (6 essays $\times$ 3 annotators). Generator-wise means summarize central tendency, while standard deviations quantify dispersion in rubric scores across prompts and grade levels.
Figure 6 provides a distributional view of total rubric outcomes, whereas Tables 6 and 7 provide complementary descriptive summaries that separate grade-level and generator-level variation. Taken together, these results indicate that rubric-guided auditing produces comparable scoring patterns across sources within an explicit analytic framework, while clarifying whether observed variation reflects writing level rather than generator identity.
B.3 Creativity Diagnostics
The LREAD rubric includes two creativity-related criteria, one under Content and one under Expression. Because creativity scores are sparse (many evaluations receive zero), we report mean $\pm$ SD and the proportion of non-zero scores (%>>0), which reflects creativity cues under the rubric.
@@ B.3 Creativity Diagnostics
Table 8 indicates higher incidence of non-zero creativity scores for human-authored essays than for most LLM sources, particularly on the content-side criterion. Some LLMs (e.g., GPT-4o and Solar) rarely trigger these creativity cues under the rubric, whereas other models exhibit a higher but still lower-than-human incidence. These patterns align with rubric-guided evaluation that rewards localized deviation and idea-level novelty beyond uniformly fluent realization.
Appendix C Taxonomy of Calibrated Human Detection Signatures
We synthesize model-wise fingerprints articulated by human annotators after completing Phase 3 and receiving gold labels for all phases. We treat these reflections as calibrated qualitative evidence of which cues became operational through iterative feedback and rubric-guided reasoning, rather than as claims about intrinsic or universal model traits. The goal is to document which linguistic regularities emerged as diagnostic for humans within our controlled essay evaluation setting.
Following the analytic framework introduced in Section 5, we organize the signatures along two axes: form-level cues—including macro-structure, connective usage, punctuation, and sentence rhythm—and content-level cues—including stance strength, evidential style, and preferred resolution patterns. This bifurcation instantiates the distinction between surface-level mechanisms and discourse-level priors that grounds the human–LLM comparison in Table 5.
C.1 Signature Elicitation Protocol
@@
Annotator notes were coded into a shared cue taxonomy through iterative consolidation. We prioritized patterns that (i) were independently reported by multiple annotators or (ii) sharply differentiates models under identical prompting and scoring regimes. Crucially, we report these signatures as after-calibration descriptors: they capture the cues that become salient once annotators are provided with gold labels and explicit rubrics, rather than impressions formed during pre-calibration or zero-shot inspection.
Model Form-Level Cues (Structure, Rhythm, Mechanics) Content-Level Cues (Stance, Evidence, Resolution)
GPT-4o Strongly standardized essay scaffold: hook $\rightarrow$ pivot $\rightarrow$ thesis in the introduction; frequent ordinal enumeration in the body; conclusions often begin with conventional discourse markers. Annotators consistently identified optional-but-recurrent comma insertion around sentence-initial connectives and following clause-linking endings. Low-risk and socially agreeable argumentation with restrained stance strength. Essays frequently close with forward-looking prescriptions and projecting generalized future outcomes, often invoking inclusive subjects (e.g., “we”) to suggest directions for society or readers. Unlike Qwen2’s emphasis on institutional duty, GPT-4o frames collective action as aspirational rather than obligatory.
Solar Connective-heavy realization with a narrow connective repertoire and repeated tokens within a single essay. Sentence rhythm tends to be short and uniform, with frequent comma insertion in positions that are grammatically permissible yet stylistically optional. Despite the frequent use of explicit connective devices, local coherence remains uneven and episodic. High surface redundancy with comparatively thin elaboration. Argument flow sometimes oscillates between claim and rebuttal moves, and explicit self-markers of opinion (e.g., “I think”, “my view”) appear performative rather than evidential.
Qwen2 Relatively natural punctuation rhythm; frequent use of formal connectives to construct explicit logical chains; frequent nominalizations and dense Sino-Korean lexicon; occasional spacing errors at multiword boundaries; recurrent quotation-like emphasis around abstract nouns. Annotators consistently identified the presence of lexical calques and rhetorical non-nativeness, localized primarily within syntactically complex modifier-noun constructions. Recurring educational or institutional framing, with conclusions emphasizing collective duty and responsibility. Figurative language appears comparatively often, while the overall register is frequently judged overly mature for the nominal student level.
@@
Llama3.1 Highest stylistic variance and frequent template mixing, weakening a single dominant structural cue. Outputs are comparatively shorter, with variable comma consistency even within the same essay. Hypothetical and counterfactual framing (“if”-clauses) appears frequently, alongside localized grammatical awkwardness. More concrete elaboration and careful counterargument handling than other models, including detailed examples. However, lexical repetition can recur without paraphrase across distant spans. A more vivid writer persona sometimes emerges through rhetorical questions or conversational devices.
@@
Table 9: After-calibration model signatures reported by human annotators following Phase 3.
C.2 Model-Specific Detection Signatures
Table 9 presents the resulting LLM-wise signature matrix for the four source LLMs (GPT-4o, Solar, Qwen2, and Llama3.1). Each row summarizes the dominant form-level and content-level cues that annotators repeatedly associated with a given LLM after calibration. The table provides a compact inventory of operationalized human diagnostics, reflecting both what annotators noticed and which cues they learned to prioritize when distinguishing LLM sources across phases and prompts.
C.3 Shared Detection Signatures
@@
Beyond model-specific fingerprints, Table 9 also reveals several cross-model regularities that annotators consistently treated as signals of machine authorship. First, LLM outputs exhibit unusually stable surface correctness relative to student writing, including fewer orthographic errors and more controlled sentence lengths. Second, outputs tend to flatten grade-level variation: essays nominally generated at elementary, middle, and high-school levels converge toward a similar band in polish and abstraction. Third, many essays converge on generic, institution-facing remedies (e.g., education, awareness campaigns, or system improvements) rather than grounded, personally constrained resolutions.
Appendix D Translationese Diagnostics in Calibrated Human Detection
Expanding on the analysis in Section 5.2, we operationalize translationese as a formal framework to account for the systematic stylistic shifts documented in Table 9, which are most pronounced in Qwen2 and notably present in GPT-4o. We further propose a set of measurement candidates that formalize this qualitative intuition as an operationalized family of cues.
D.1 Motivation and Prior Evidence
We use translationese to denote systematic source-language fingerprints and stylistic shifts that distinguish translated text from natively authored text. Such patterns can be learned and propagated during model pretraining when translationese is present in the training data (doshi-etal-2024-pretraining). Multilingual evaluation pipelines that rely on translated instructions may further confound judgments of naturalness and bias comparisons between English-conditioned and target-language behaviors (enomoto-etal-2025-fair). In multilingual LLMs, supervised training objectives can incentivize overly literal realizations that preserve semantic adequacy while degrading native-like naturalness (li-etal-2025-lost). Relatedly, prior work reports an “English accent” effect in multilingual generation, which reduces perceived target-language naturalness (guo-etal-2025-large).
D.2 Translationese Diagnostics
@@
Grounded in the calibrated human signature inventory (Table 9), we outline a set of translationese-sensitive diagnostics that can be instantiated as measurable features:
@@
  • Nominalization Rate: the proportion of nominalized predicates and abstract nominal constructions, capturing “noun-heavy” discourse packaging characteristic of literal transfer.
@@
  • Discourse Deixis and Pronominal Reference: frequencies of demonstratives and pronominal forms (e.g., this/that-type anaphora) that are associated with English-style translated discourse structuring, particularly salient in GPT-4o outputs.
@@
  • Sentence-Level Naturalness Scoring: complementary global measures of perceived naturalness, following prior work on multilingual naturalness evaluation (guo-etal-2025-large).
These diagnostics are designed to shift translationese from an impressionistic descriptor to a reproducible family of cues. They make explicit how calibrated human detection can leverage language-specific micro-realization patterns not captured by coarse fluency or document-level coherence.

### v2

@@ B.4 Statistical Tests
Test $p$-value / effect
Phase 1 majority vote vs. chance (50%) Binomial, $p=0.181$
[…]
False-negative reduction on AI essays Fisher’s exact, $p=0.003$
Human vs. LLM-judge majority vote (Phase 1) McNemar exact, $p<0.001$
Table 9: Inferential tests corresponding to the main claims in the paper.
B.5 Rubric-Score Diagnostics
Figure 6 and Tables 1012 provide a graded view of rubric-guided judgments beyond binary attribution. Across pooled evaluations, LLM essays tend to score higher on formal expression-related criteria, whereas human essays more often receive non-zero creativity scores. This inversion helps explain why intuition-only readers can be misled by surface correctness and why rubric-guided evaluation improves discrimination.
Figure 6: Distribution of mean Phase 2 LREAD scores by source group, aggregated over three annotators for each essay. Higher scores indicate stronger formal conformity to rubric criteria, not necessarily stronger evidence of human authorship.
@@ B.5 Rubric-Score Diagnostics
Table 10: Phase 2 total LREAD scores by grade level (mean $\pm$ SD). Scores are summarized for each annotator and for pooled evaluator–essay evaluations. Each grade level contains 10 essays, yielding 30 pooled evaluations per level across three annotators.
@@
Appendix C Statistical Reporting, Voting Conventions, and Prompt Protocols
This appendix documents implementation details that are important for reproducibility: voting conventions, uncertainty handling, statistical procedures, and prompt protocols for both human and LLM-based reviewing.
C.1 Voting Conventions
We use O = AI, X = Human, and $\triangle$ = uncertain throughout the annotation protocol. Individual annotators are allowed to use $\triangle$ when they judge a case to be insufficiently resolved by the available evidence.
For individual-level accuracy, $\triangle$ is counted as incorrect. This is consistent with the task objective: the blind evaluation requires a binary Human-versus-AI decision, and uncertainty does not itself constitute a correct attribution.
For Fleiss’ $\kappa$, $\triangle$ is treated as a third category. We adopt this inclusive formulation because uncertainty is a meaningful part of the calibration process, especially in earlier phases, and excluding such responses would artificially remove difficult cases from the agreement analysis.
For majority vote, $\triangle$ counts as one vote. When all three annotators choose different categories (O / X / $\triangle$), we conservatively resolve the essay to Human (X) in order to minimize false accusation. Under this rule, uncertain responses never override a binary majority.
C.2 Statistical Procedures
Accuracy confidence intervals are reported using the exact Clopper–Pearson method. We use exact intervals because the sample sizes are small, especially in Phase 3.
Confidence intervals for Fleiss’ $\kappa$ are estimated by nonparametric bootstrap over items, with item-level resampling and B=20, [^0] B=20{,}000 bootstrap samples. We report percentile intervals. Because Phase 3 contains only 10 essays, $\kappa$ intervals in that phase are unstable and are therefore interpreted cautiously alongside unanimity counts.
[…]
For the comparison between the human majority vote and the LLM-judge majority vote on the same Phase 1 essays, we use the exact McNemar test. This comparison is appropriate because the two majority-vote systems are applied to the same underlying item set.
C.3 Human Reflection Procedure After Phase 1
After the Phase 1 blind task and gold-label disclosure, each annotator wrote short structured reflections describing (i) which cues they had relied on, (ii) which essays they had found difficult, (iii) which initial assumptions appeared unreliable after gold-label disclosure, and (iv) which recurrent artifacts seemed potentially source-specific or human-typical. These reflections were used as inputs to the collaborative Phase 2 rubric-construction process.
The purpose of this reflection step was not to optimize performance on the Phase 1 items, but to externalize previously implicit heuristics so that they could later be re-anchored to explicit scoring criteria.
C.4 Human Signature Elicitation After Phase 3
After completing the blind Phase 3 task and receiving gold labels, annotators wrote structured notes on recurring model-specific patterns and on cues associated with authentic human writing. These notes included both form-level observations (e.g., structure, punctuation, spacing, connective usage, sentence rhythm) and content-level observations (e.g., stance, evidential style, resolution patterns, persona consistency).
The resulting notes were consolidated through iterative discussion into the shared cue taxonomy reported in Appendix D. We treat these signatures as after-calibration descriptors, not as universal or immutable properties of the underlying models.
C.5 LLM-Judge Blind Attribution Prompt
For the zero-shot LLM-reviewer baseline, each model first receives a blind binary-attribution instruction. The English translation of the prompt is as follows:
“You are a reviewer who determines the author of a Korean essay. The essay is either written by a human or generated by an AI model. The writer is an elementary-, middle-, or high-school student. If the essay is AI-generated, the AI produced it at an elementary-, middle-, or high-school level. Decide whether the author is human or AI. Answer X for human and O for AI.”
This prompt is intentionally minimal. Its purpose is to estimate practical zero-shot reviewing behavior rather than optimized prompted performance.
C.6 LLM-Judge Post-hoc Cue Elicitation Prompt
After completing the blind O/X decision task, each LLM reviewer is shown the gold labels and generator identities and is then asked to summarize the characteristics it noticed for each source. The English translation of the elicitation prompt is as follows:
“Summarize the characteristics you noticed for each author type during your review. Organize the summary for Human, GPT-4o, Solar, Qwen2, and Llama3.1. Use the format <Form> and <Content>.”
We use this second-stage prompt only for qualitative comparison with the human cue inventory. It is not part of the blind detection decision itself.
C.7 Runtime Conditions for LLM Judges
LLM judges were run through provider-facing web interfaces using provider-default decoding settings. We did not manually tune temperature, top-$p$, or other generation controls. We report this setup to reflect realistic zero-shot reviewing conditions rather than a carefully optimized prompting pipeline.
The judge models are used as practical reviewing baselines, not as calibrated or task-specialized detectors. Their role in the paper is therefore contrastive: they provide a useful benchmark for what strong zero-shot reviewing can achieve, while LREAD provides an auditable framework for explicit human reasoning.
C.8 Original Korean Prompt Text
Blind attribution prompt.
너는 지금부터 한국어로 작성된 에세이의 작성자를 판별하는 검수자야. 내가 제공하는 에세이는 사람이 작성한 글이거나 AI 모델이 생성한 글이야. 이때 글을 작성한 사람은 초등학생이거나 중학생이거나 고등학생이야. AI 모델이 생성한 경우에 AI 모델은 초등학생 수준 혹은 중학생 수준 혹은 고등학생 수준으로 글을 생성했어. 내가 제공하는 글을 쓴 주체가 사람인지 AI인지 판단해줘. 사람이면 X로 대답하고 AI면 O로 대답해줘.
Post-hoc cue elicitation prompt.
너가 검수를 하면서 느낀 작성자별 특징을 정리해줘. 사람, GPT-4o, Solar, Qwen2, Llama3.1 각각 특징을 정리해줘. 아래 양식에 맞춰서 작성해줘. <글 형식> <글 내용>
Appendix D Calibrated Cue Taxonomy and Model-Specific Signatures
This section documents the after-calibration cue inventory articulated by human annotators. We treat these cues as calibrated descriptors that became operational within our controlled study, not as universal or immutable properties of specific models. The purpose of this appendix is to show how rubric-guided calibration transforms diffuse impressions into an inspectable signature inventory that can be reviewed, debated, and adapted.
D.1 Signature Elicitation Protocol
After completing the blind Phase 3 task, annotators reviewed gold labels and wrote structured reflections on (i) recurring source-specific patterns, (ii) cues associated with authentic human writing, and (iii) disagreement-triggering cases from earlier phases. These notes were then consolidated into a shared cue taxonomy through iterative discussion.
@@ D.1 Signature Elicitation Protocol
We prioritized patterns that were either independently reported by multiple annotators or that sharply differentiated sources under identical prompting and rubric conditions. We report these signatures as after-calibration descriptors: they capture cues that became operational once annotators had explicit criteria and feedback, rather than impressions formed during zero-shot inspection.
D.2 Model-Specific Signatures
Across annotators, authentic human writing was associated with localized developmental errors, uneven but grounded elaboration, and persona-consistent everyday constraints. By contrast, the model-specific patterns below summarize the recurring cues attributed to each LLM family after calibration. Table 13 should be read as an auditable inventory of learned human diagnostics, not as a permanent taxonomy of model essence.
Generator Form-level cues Content-level cues Persona / affect cues Caution / overlap note
GPT-4o standardized essay scaffold; ordinal transitions; polished sentence boundaries; optional connective-comma insertion low-risk socially agreeable claims; forward-looking prescriptions; balanced concession moves inclusive “we” voice; mature but restrained stance; weak local lived detail can overlap with carefully coached school essays; diagnose jointly with persona fidelity and punctuation micro-patterns
Solar connective-heavy sentences; narrow connective repertoire; short rhythmic clauses; repeated optional commas redundancy with thin elaboration; claim–rebuttal oscillation; performative opinion markers emotionally assertive but shallow; less stable persona control distinguish from formulaic human writing by repetition density and weak evidential progression
Qwen2 formal connective chains; nominalization-heavy diction; dense Sino-Korean lexicon; occasional spacing errors; quotation-like emphasis educational / institutional framing; collective-duty arguments; abstract justification overly mature, adult-like register for student persona; non-native or calqued phrasing strongest translationese-sensitive profile; evaluate jointly with natural-Korean-phrasing criteria
@@ D.2 Model-Specific Signatures
Llama3.1 highest stylistic variance; template mixing; shorter outputs; variable comma consistency; occasional grammatical awkwardness more concrete elaboration; careful counterarguments; detailed examples; lexical repetition without paraphrase comparatively vivid persona through rhetorical questions or conversational devices can appear more human-like than other generators; use multiple cues rather than any single marker
@@
Table 13: Calibrated model-specific signatures elicited from human annotators after Phase 3. These are after-calibration descriptors, not immutable properties of the models.
D.3 Shared Detection Signatures
@@
Beyond model-specific fingerprints, the shared cue inventory reveals several cross-model regularities that annotators consistently treated as signals of machine authorship. First, LLM outputs exhibit unusually stable surface correctness relative to student writing, including fewer orthographic errors and more controlled sentence lengths. Second, outputs tend to flatten grade-level variation: essays nominally generated at elementary, middle, and high-school levels converge toward a similar band of polish and abstraction. Third, many essays converge on generic, institution-facing remedies (e.g., education, awareness campaigns, or system improvements) rather than grounded, personally constrained resolutions.
These regularities should be interpreted as family-level hypotheses that motivate adaptation and revalidation, not as permanent signatures guaranteed to survive model updates.
Appendix E Prospective Translationese Diagnostics for Future Operationalization
Building on the calibrated cue inventory in Appendix D, we outline translationese-sensitive diagnostics that may be operationalized in future work. We do not compute these measures in the present study; instead, we include them to clarify how an impressionistic human cue can be translated into a reproducible measurement agenda.
The proposed diagnostics correspond primarily to the Expression dimension of LREAD, especially natural Korean phrasing, register stability, and connective / punctuation behavior.
E.1 Motivation and Prior Evidence
We use translationese to denote systematic source-language fingerprints and stylistic shifts that distinguish translated text from natively authored text. Prior work suggests that such patterns can be learned and propagated during model pretraining, amplified by multilingual instruction pipelines, and reflected in reduced target-language naturalness. The cue inventory in Appendix D indicates that translationese-sensitive distortions are especially salient for Qwen2 and, to a lesser degree, GPT-4o in our dataset.
E.2 Translationese Diagnostics
@@
Grounded in the calibrated human signature inventory, we outline a set of translationese-sensitive diagnostics that can be instantiated as measurable features:
@@
  • Nominalization Rate: the proportion of nominalized predicates and abstract nominal constructions, capturing noun-heavy discourse packaging characteristic of literal transfer.
@@
  • Discourse Deixis and Pronominal Reference: frequencies of demonstratives and pronominal forms (e.g., this/that-type anaphora) that are associated with English-style translated discourse structuring.
@@
  • Sentence-Level Naturalness Scoring: complementary global measures of perceived naturalness, following prior work on multilingual naturalness evaluation.
These diagnostics are designed to shift translationese from an impressionistic descriptor to a reproducible family of cues. The goal is not to claim that translationese has already been quantitatively validated here, but to document a plausible path from calibrated human perception to future operational measurement.